# An embedding-based distance for temporal graphs

### v2

@@ Introduction
Figure 3: Distance-based graph clustering for ensembles of temporal graphs generated according to different randomization techniques. Left panel: results for the matched distance $d_{\rm m}$ of (3). Right panel: results for the unmatched distance $d_{\rm u}$ of (4). Within each panel, each matrix corresponds to one of the $9$ SocioPatterns temporal graphs described in Table 1. The rows and columns of each matrix correspond to the same set of $6$ randomization techniques we used, described in the Methods sections. Each input graph is represented as a sequence of temporal edges $(i,j,t)$ as per Definition 1. The randomizations act on the temporal edges as follows. Random: preserves the number of temporal edges and randomizes the node and time indices; Random delta: preserves the number of temporal edges and interaction duration distribution; Active snapshot: preserves the number of edges at each time-step and the times at which each node is active, i.e., at which it has at least one neighbor. Time: preserves the aggregated weighted graph structure, i.e., the number of times each edge is active in the temporal graph; Sequence; preserves each snapshot’s adjacency matrix and randomizes the order in which they appear; Weighted degree: preserves the total number of temporal edges involving each node. For each pair of randomizations, we infer the randomization method of each temporal graph via an unsupervised distance-based clustering algorithm, and we compare the inferred randomization method with the known true one. Each matrix entry reports (value and color coding) the accuracy of the inferred labels (randomization methods), quantified as the normalized mutual information (NMI) between the inferred and true labels.
@@
To quantify the discriminative performance of our temporal graph distance, we proceed as follows, separately for each empirical temporal graph: given the empirical temporal graph and a pair of randomization operations ($R_{i}$, $R_{j}$), we generate a set of $250$ realizations using $R_{i}$, labeled as 00, and a set of $250$ realizations using $R_{j}$, labeled as $1$, for a total of $500$ temporal graphs. We then compute the matrix of distances between these $500$ graphs and use it to cluster them in $2$ clusters, following the same procedure used for Fig. 2A. Figure 3 reports the NMI values between the real labels and the unsupervised algorithm result for each pair of randomizations ($R_{i}$, $R_{j}$), and for both the matched distance $d_{\rm m}$ (left panel) and the unmatched distance $d_{\rm u}$ (right panel). Within each panel, each matrix refers to one empirical temporal graph, and its entries are the NMI values described above. A high NMI entry indicates that our distance discriminates well the temporal graphs generated according to the two randomization procedures in the corresponding row and column of the matrix entry, i.e., that the distance is sensitive to the properties preserved by either $R_{i}$ or $R_{j}$. Low NMI values may either correspond to an inability of the distance to capture those properties or simply reflect that the empirical temporal graph is statistically random as far as those properties are concerned, as observed in particular for several pairs of randomizations in the Baboons data set. Thus, it is enough to find one empirical data set for which the distance can distinguish between a pair of randomizations to conclude that the distance is indeed sensitive to the corresponding graph properties.
Conclusions
@@
We introduced a novel definition of distance between pairs of temporal graphs. This definition entails two steps. First, the temporal graphs are embedded in Euclidean space, and then a distance is defined in embedded space. For the first step, we use an embedding based on time-respecting random walks over the temporal graph. Such walks are known to depend on and to encode important structural and temporal properties of time-varying graphs [70, 9, 58, 43, 59]. For the second step, we proposed two possibilities for the distance definition: if a mapping is known between the nodes of the graphs to be compared, we consider a distance definition that leverages such mapping; in the more general case, when such a mapping is unavailable, we put forward a definition that makes it possible to compare graphs with arbitrarily different sizes (numbers of nodes). In both cases, since the size of the embedding matrix we use does not depend on the graph’s temporal span, it is possible to embed temporal graphs with different durations in the same embedding space and thus compute a distance between them. The evaluation of our approach on both synthetic and empirical data shows that the proposed distance is sensitive to structural differences (e.g., degree distribution, clustering coefficient, or presence of communities) as well as to temporal differences (e.g., burstiness or node activity patterns) of the temporal graphs being compared.
@@
Defining a distance between temporal graphs opens the door to a wealth of potential applications, of which our evaluation only offers possible examples. On the one hand, such a distance can support data analysis of temporal graph data by extending the procedure of Figure 3 to a richer set of randomizations [50]. Indeed, as discussed, the inability to distinguish pairs of randomizations of a given temporal graph can be regarded as an invariance property of the graph itself. On the other hand, computing a distance between graphs can provide a crucial tool for the validation of generative models: the validation of generative models is often limited to comparing a set of statistical properties [41], while our distance definition enables us to carry out a comparison at the global level. In addition, our distance could underpin the very process of generating synthetic temporal graphs by enabling approaches such as GANS [71] that depend on a global distance and have proved extremely powerful in generating realistic data [72, 73]. In particular, generating synthetic human proximity data with realistic topological and temporal properties could be used to better simulate infectious disease spread and other phenomena of interest for public health research [74, 75, 76, 77, 78, 79]. Finally, we know that empirical (temporal) graphs are hard to anonymize [80, 81, 78]. To minimize the risk of node re-identification, performing some perturbation operations on the graph might be necessary before making it publicly available. However, such operations risk destroying essential patterns and information of empirical data. The distance we introduced could help tackle the trade-offs between re-identification risk and information loss by better quantifying the latter.
@@ Methods
The EDRep algorithm [60] was recently proposed to efficiently generate embeddings given a probability matrix encoding the affinity between the embedded items. This is done by optimizing the cost function of (2) under the constraint $\|\bm{x}_{i}\|=1$ and obtaining a low-dimensional representation of the information encoded in the matrix $P$. A known problem of this type of cost function is the computational complexity because the normalization constants $Z_{i}=\sum_{k\in\mathcal{V}}e^{\bm{x}_{i}^{T}\bm{x}_{k}}$ require $\mathcal{O}(n^{2})$ to be calculated. However, Ref. [60] describes an efficient way to estimate all the $Z_{i}$ values in $\mathcal{O}(n)$. This is accomplished by first subdividing the nodes into $q$ groups based on the embedding matrix $X$. Here $q$ is a parameter of the algorithm, and larger $q$ values generally lead to a higher accuracy, but very good results are obtained already for $q=1$. Second, for each $a=1,\dots,q$, one computes the mean $\bm{\mu}_{a}$ and the covariance matrix $\Omega_{a}$ of the set of embedding vectors of the nodes in group $a$. Denoting the number of nodes in group $a$ by $\pi_{a}$, one obtains the estimation of $Z_{i}$ as
$Z_{i}\approx\sum_{a=1}^{q}\pi_{a}~{}{\rm exp}\left\{\bm{x}_{i}^{T}\bm{\mu_{a}}+\frac{1}{2}\bm{x}_{i}^{T}\Omega_{a}\bm{x}_{i}\right\}\,\,.$
Coming to the algorithm’s computational complexity, given that our choice of $P$ is the sum of matrix products, the algorithm can be deployed in two forms: one in which $P$ is never formally computed and the sequence $\{\hat{L}_{t}\}_{t=1,\dots,T}$ is fed to the algorithm; and one in which $P$ is explicitly computed. Letting $X\in\mathbb{R}^{n\times d}$ be the matrix storing the embedding vectors in its rows, the computational complexity is determined by the matrix product $PX$ needed to obtain the gradient of $\mathcal{L}$. In the former case, this is computed in $\mathcal{O}(nd^{2}+dE)$ operations, where $E=\sum_{t=1}^{T}|\mathcal{E}_{t}|$ is the number of temporal edges. In the second case, instead, the complexity is given by the number of non-zero entries of $P$, which cannot exceed $n^{2}$. The former implementation is particularly convenient when $n$ is large, and the $\hat{L}_{t}$ matrices are very sparse. This is because $P$ may be dense, even if the snapshots are sparse, and for large $n$, $P$ might not even fit in the computer’s memory. Conversely, for small graphs, $n^{2}$ may be smaller than the number of temporal edges, making the latter implementation more convenient.
@@
$\bar{X}_{ia}:=\left(QX\right)_{ia}=\sum_{j\in\mathcal{V}}Q_{ij}X_{ja}=\sum_{j\in\mathcal{V}}\delta_{j,\pi(i)}X_{ja}=X_{\pi(i),a}.$
@@
$(QQ^{T})_{ij}=\sum_{k\in\mathcal{V}}Q_{ik}Q_{jk}=\sum_{k\in\mathcal{V}}\delta_{k,\pi(i)}\delta_{k,\pi(j)}=\delta_{\pi(i),\pi(j)}=\delta_{ij},$
@@
$\bar{X}^{T}\bar{X}=X^{T}\underbrace{Q^{T}Q}_{I_{n}}X=X^{T}X.$
We now show that the distance $d_{\rm u}$ is also invariant under the orthogonal transformation applied to the embedding matrix rows. Let $R\in\mathbb{R}^{d\times d}$ be an orthogonal matrix and let $\tilde{X}=XR$. Letting $\lambda_{i}(M)$ be the $i$-th smallest eigenvalue of $M$, then, for all $i$, $\lambda_{i}(AB)=\lambda_{i}(BA)$ [82, Theorem 1.3.22]. It follows
$\lambda_{i}(\tilde{X}^{T}\tilde{X})=\lambda_{i}(R^{T}X^{T}XR)=\lambda_{i}(\underbrace{RR^{T}}_{I_{d}}X^{T}X)=\lambda_{i}(X^{T}X),$
@@
In Table 1 we summarize the properties of the temporal graphs we used to conduct our tests.
We now give a more detailed description of the randomization techniques we adopted, defined in terms of the quantities they preserve. According to the method, it may be more convenient to represent the temporal graph as a sequence of instantaneous interactions $(i,j,t)$, as a sequence of interactions with a duration $(i,j,t,\tau)$, or as a sequence of weighted adjacency matrices $W^{(t)}$ [69]. Before randomization, time is discretized at the scale of $10$ minutes in each data set, and the cumulative interaction time in each $10$-minute window is used as weight.
@@
    Random. Temporal edges are represented as $(i,j,t)$ and all three indices are randomized, avoiding self-edges ($i$ and $j$ randomized between $1$ and $n$, and $t$ between $1$ and $T$).
@@
    Random delta. Temporal edges are represented as $(i,j,t,\tau)$. Once again, $i,j,t$ are randomized ($t$ is sampled between $1$ and $T-\tau$), while $\tau$ is preserved.
@@
    Time. Temporal edges are represented as $(i,j,t)$ and only the index $t$ is randomized.
@@
    Sequence. The graph is represented as a sequence of weighted adjacency matrices $W^{(t)}$ and the randomization acts on the indices $t$.
@@
Table 1: Summary properties of the SocioPatterns time-resolved proximity networks used here. Graph name is used to identify the graphs; Description provides concise information on the context where data were collected; $n$ is the number of graph nodes; Duration is the temporal span of the data set, and $T$ is the number of graph snapshots.
@@
We here provide a formal definition of the models used to generate the synthetic graphs under analysis. Even though we considered four models, three of these can be generated from the degree corrected stochastic block model [63] by changing its parameters.
Definition 3 (Degree corrected stochastic block model)
Let $\mathcal{V}$ be a set of $n$ nodes and $\bm{\ell}\in[k]^{n}$ be a vector mapping each node to a class. Further let $C\in\mathbb{R}^{k\times k}$ be a positive symmetric matrix and $\bm{\theta}\in\mathbb{R}^{n}$ be a vector satisfying $\bm{\theta}^{T}\bm{1}_{n}=n$. The entries of the graph adjacency matrix $A\in\ [0,1]^{n\times n}$ are generated independently (up to symmetry) at random with probability
$\mathbb{P}(A_{ij}=1)={\rm min}\left(1,\theta_{i}\theta_{j}\frac{C_{\ell_{i},\ell_{j}}}{n}\right)\,.$
The vector $\bm{\ell}$ contains the labels and gives a community structure in the case in which $C_{a,a}>C_{a,b}$ for $b\neq a$, meaning that there is a higher probability that two nodes in the same community will get connected. The value $\theta_{i}$ is proportional to the expected degree of node $i$. For this reason, if one chooses $\bm{\theta}=\bm{1}_{n}$ and $\bm{\ell}=\bm{1}_{n}$, one gets the Erdős-Renyi model, in which every node has the same expected degree, and there are no communities. The configuration model, instead, is obtained by letting $\bm{\ell}=\bm{1}_{n}$, but changing the value of $\bm{\theta}$ to create an arbitrary degree distribution that we choose to be a (properly rescaled) uniform distribution between $3$ and $10$ raised to the power $4$. Finally, the stochastic block model is obtained from a labeling vector different from $\bm{1}_{n}$ and letting $\bm{\theta}=\bm{1}_{n}$. We consider $k=5$ communities of equal size with $C_{a,b}=20\delta_{ab}+(1-\delta_{ab})$, with $\delta$ the Kroeneker symbol.
@@
Let $\mathcal{V}$ be a set of $n$ nodes. For each $i\in\mathcal{V}$ let $\bm{x}_{i}\in\mathbb{R}^{2}$ be a random vector with norm $\|\bm{x}_{i}\|\leq 1$. The entries of the graph adjacency matrix $A\in[0,1]^{n\times n}$ are generated independently (up to symmetry) at random with probability
$\mathbb{P}(A_{ij}=1)=e^{-\beta\|\bm{x}_{i}-\bm{x}_{j}\|}\,,$
@@
Note that even though the entries of $A$ are drawn at random, this model can generate graphs with a high clustering coefficient because the probability depends on the relative distance of a fixed embedding. In our simulations, we set $\beta=20$.

### v3

@@ Results
Figure 4: Distance-based graph clustering for ensembles of temporal graphs generated according to different randomization techniques. Left panel: results for the matched distance $d_{\rm m}$ of Eq. (5). Right panel: results for the unmatched distance $d_{\rm u}$ of Eq. (6). Within each panel, each matrix corresponds to one of the $9$ SocioPatterns temporal graphs described in Table 1. The rows and columns of each matrix correspond to the same set of $6$ randomization techniques we used, described in the Methods sections. Each input graph is represented as a sequence of temporal edges $(i,j,t)$ as per Definition 1. The randomizations act on the temporal edges as follows. Random: preserves the number of temporal edges and randomizes the node and time indices; Random delta: preserves the number of temporal edges and interaction duration distribution; Active snapshot: preserves the number of edges at each time-step and the times at which each node is active, i.e., at which it has at least one neighbor. Time: preserves the aggregated weighted graph structure, i.e., the number of times each edge is active in the temporal graph; Sequence; preserves each snapshot’s adjacency matrix and randomizes the order in which they appear; Weighted degree: preserves the total number of temporal edges involving each node. For each pair of randomizations, we infer the randomization method of each temporal graph via an unsupervised distance-based clustering algorithm, and we compare the inferred randomization method with the known true one. Each matrix entry reports (value and color coding) the accuracy of the inferred labels (randomization methods), quantified as the normalized mutual information (NMI) between the inferred and true labels.
@@
To quantify the discriminative performance of our temporal graph distance, we proceed as follows, separately for each empirical temporal graph: given the empirical temporal graph and a pair of randomization operations ($R_{i}$, $R_{j}$), we generate a set of $250$ realizations using $R_{i}$, labeled as 00, and a set of $250$ realizations using $R_{j}$, labeled as $1$, for a total of $500$ temporal graphs. We then compute the matrix of distances between these $500$ graphs and use it to cluster them in $2$ clusters, following the same procedure used for Fig. 2A. Figure 4 reports the NMI values between the real labels and the unsupervised algorithm result for each pair of randomizations ($R_{i}$, $R_{j}$), and for both the matched distance $d_{\rm m}$ (left panel) and the unmatched distance $d_{\rm u}$ (right panel). Within each panel, each matrix refers to one empirical temporal graph, and its entries are the NMI values described above. A high NMI entry indicates that our distance discriminates well between the temporal graphs generated according to the two randomization procedures in the corresponding row and column of the matrix entry, i.e., that the distance is sensitive to the properties preserved by either $R_{i}$ or $R_{j}$. Low NMI values may either correspond to an inability of the distance to capture those properties or simply reflect that the empirical temporal graph is statistically random as far as those properties are concerned, as observed in particular for several pairs of randomizations in the Baboons dataset. Thus, it is enough to find one empirical dataset for which the distance can distinguish between a pair of randomizations to conclude that the distance is indeed sensitive to the corresponding graph properties.
Discussion
@@
We introduced a novel definition of distance between pairs of temporal graphs. This definition entails two steps. First, the temporal graphs are embedded in Euclidean space, and then a distance is defined in embedding space. For the first step, we use an embedding based on time-respecting random walks over the temporal graph. Such walks are known to depend on and to encode important structural and temporal properties of time-varying graphs [71, 9, 59, 43, 60]. For the second step, we proposed two possibilities for the distance definition: if a mapping is known between the nodes of the graphs to be compared, we consider a distance definition that leverages such mapping; in the more general case, when such a mapping is unavailable, we put forward a definition that makes it possible to compare graphs with arbitrarily different sizes (numbers of nodes). In both cases, since the size of the embedding matrix we use does not depend on the graph’s temporal span, it is possible to embed temporal graphs with different durations in the same embedding space and thus compute a distance between them. The evaluation of our approach on both synthetic and empirical data shows that the proposed distance is sensitive to structural differences (e.g., degree distribution, clustering coefficient, or presence of communities) as well as to temporal differences (e.g., burstiness or node activity patterns) of the temporal graphs being compared.
@@
Defining a distance between temporal graphs opens the door to a wealth of potential applications, of which our evaluation only offers possible examples. On the one hand, such a distance can support data analysis of temporal graph data by extending the procedure of Figure 4 to a richer set of randomizations [51]. Indeed, as discussed, the inability to distinguish pairs of randomizations of a given temporal graph can be regarded as an invariance property of the graph itself. On the other hand, computing a distance between graphs can provide a crucial tool for the validation of generative models that is often limited to comparing a set of statistical properties [41], while our distance definition enables us to carry out a comparison at the global level. In addition, our distance could underpin the very process of generating synthetic temporal graphs by enabling approaches such as GANS [72] that depend on a global distance and have proved extremely powerful in generating realistic data [73, 74]. In particular, generating synthetic human proximity data with realistic topological and temporal properties could be used to better simulate infectious disease spread and other phenomena of interest for public health research [75, 76, 77, 78, 79, 80]. Finally, we know that empirical (temporal) graphs are hard to anonymize [81, 82, 79]. To minimize the risk of node re-identification, performing some perturbation operations on the graph might be necessary before making it publicly available. However, such operations risk destroying essential patterns and information of empirical data. The distance we introduced could help tackle the trade-offs between re-identification risk and information loss by better quantifying the latter.
@@ Methods
The EDRep algorithm [61] was recently proposed to efficiently generate embeddings given a probability matrix encoding the affinity between the embedded items. This is done by optimizing the cost function of Eq. (2) under the constraint $\|\mathbf{x}_{i}\|=1$ and obtaining a low-dimensional representation of matrix $P$. A known problem of this type of cost function is the computational complexity because the normalization constants $Z_{i}=\sum_{k\in\mathcal{V}}e^{\mathbf{x}_{i}^{T}\mathbf{x}_{k}}$ require $\mathcal{O}(n^{2})$ to be calculated. However, Ref. [61] describes an efficient way to estimate all the $Z_{i}$ values in $\mathcal{O}(n)$. This is accomplished by first subdividing the nodes into $q$ groups based on the embedding matrix $X$. Here $q$ is a parameter of the algorithm, and larger $q$ values generally lead to a higher accuracy, but very good results are already obtained for $q=1$. Second, for each $a=1,\dots,q$, one computes the mean $\bm{\mu}_{a}$ and the covariance matrix $\Omega_{a}$ of the set of embedding vectors of the nodes in group $a$. Denoting the number of nodes in group $a$ by $\pi_{a}$, one obtains the estimation of $Z_{i}$ as
$Z_{i}\approx\sum_{a=1}^{q}\pi_{a}~{}{\rm exp}\left\{\mathbf{x}_{i}^{T}\bm{\mu}_{a}+\frac{1}{2}\mathbf{x}_{i}^{T}\Omega_{a}\mathbf{x}_{i}\right\}\,\,.$ (7)
Coming to the algorithm’s computational complexity, given that our choice of $P$ is the sum of matrix products, the algorithm can be implemented in two forms: one in which $P$ is never formally computed and the sequence $\{\hat{L}_{t}\}_{t=1,\dots,T}$ is fed to the algorithm; and one in which $P$ is explicitly computed. Letting $X\in\mathbb{R}^{n\times d}$ be the matrix storing the embedding vectors in its rows, the computational complexity is determined by the matrix product $PX$ needed to obtain the gradient of $\mathcal{L}_{P}$ defined in Eq. (2). In the former case, this is computed in $\mathcal{O}(nd^{2}+dE)$ operations, where $E=\sum_{t=1}^{T}|\mathcal{E}_{t}|$ is the number of temporal edges. In the second case, instead, the complexity is given by the number of non-zero entries of $P$, which cannot exceed $n^{2}$. The former implementation is particularly convenient when $n$ is large and the $\hat{L}_{t}$ matrices are very sparse. This happens because $P$ may be dense, even if the snapshots are sparse, and for large values of $n$, representing the matrix $P$ in computer memory might be challenging. Conversely, for small graphs, $n^{2}$ may be smaller than the number of temporal edges, making the latter implementation more convenient.
@@
$\bar{X}_{ia}:=\left(QX\right)_{ia}=\sum_{j\in\mathcal{V}}Q_{ij}X_{ja}=\sum_{j\in\mathcal{V}}\delta_{j,\pi(i)}X_{ja}=X_{\pi(i),a}\,\,.$ (9)
@@
$(QQ^{T})_{ij}=\sum_{k\in\mathcal{V}}Q_{ik}Q_{jk}=\sum_{k\in\mathcal{V}}\delta_{k,\pi(i)}\delta_{k,\pi(j)}=\delta_{\pi(i),\pi(j)}=\delta_{ij}\,\,,$ (10)
@@
$\bar{X}^{T}\bar{X}=X^{T}\underbrace{Q^{T}Q}_{I_{n}}X=X^{T}X\,\,.$ (11)
We now show that the distance $d_{\rm u}$ is also invariant under the orthogonal transformation applied to the embedding matrix rows. Let $R\in\mathbb{R}^{d\times d}$ be an orthogonal matrix and let $\tilde{X}=XR$. Letting $\lambda_{i}(M)$ be the $i$-th smallest eigenvalue of $M$, then, for all $i$, $\lambda_{i}(AB)=\lambda_{i}(BA)$ [83, Theorem 1.3.22]. It follows
$\lambda_{i}(\tilde{X}^{T}\tilde{X})=\lambda_{i}(R^{T}X^{T}XR)=\lambda_{i}(\underbrace{RR^{T}}_{I_{d}}X^{T}X)=\lambda_{i}(X^{T}X)\,\,,$ (12)
@@
In Table 1 we summarize the properties of the temporal graphs we used to conduct our tests. Here we give a more detailed description of the randomization techniques we adopted, defined in terms of the quantities they preserve. According to the method, it may be more convenient to represent the temporal graph as a sequence of instantaneous interactions $(i,j,t)$, as a sequence of interactions with a duration $(i,j,t,\tau)$, or as a sequence of weighted adjacency matrices $W^{(t)}$ [70]. Before randomization, time is discretized at the scale of $10$ minutes in each dataset, and the cumulative interaction time in each $10$-minute window is used as weight.
@@
    Random. Temporal edges are represented as $(i,j,t)$ and all three indices are randomized, avoiding self-edges ($i$ and $j$ randomly sampled with replacement between $1$ and $n$, and $t$ between $1$ and $T$).
@@
    Random delta. Temporal edges are represented as $(i,j,t,\tau)$. Once again, $i,j,t$ are randomized ($t$ is randomly sampled with replacement between $1$ and $T-\tau$), while $\tau$ is preserved.
@@
    Time. Temporal edges are represented as $(i,j,t)$ and only the index $t$ is randomly sampled with replacement.
@@
    Sequence. The graph is represented as a sequence of weighted adjacency matrices $W^{(t)}$ and the indices $t$ are shuffled.
@@
Table 1: Summary properties of the SocioPatterns time-resolved proximity networks used here. Graph name is used to identify the graphs; Description provides concise information on the context where data were collected; $n$ is the number of graph nodes; Duration is the temporal span of the dataset, and $T$ is the number of graph snapshots. The temporal resolution is $t_{\rm res}=10$ min for all datasets.
@@
We here provide a formal definition of the models used to generate the synthetic graphs under analysis. Even though we considered four models, three of these can be generated from the degree-corrected stochastic block model [64] by changing its parameters.
Definition 3 (Degree-corrected stochastic block model)
Let $\mathcal{V}$ be a set of $n$ nodes and $\bm{\ell}\in[k]^{n}$ be a vector mapping each node to a class. Further let $C\in\mathbb{R}^{k\times k}$ be a positive symmetric matrix and $\bm{\theta}\in\mathbb{R}^{n}$ be a vector satisfying $\bm{\theta}^{T}\mathbf{1}_{n}=n$. The entries of the graph adjacency matrix $A\in\ [0,1]^{n\times n}$ are generated independently (up to symmetry) at random with probability
$\mathbb{P}(A_{ij}=1)={\rm min}\left(1,\theta_{i}\theta_{j}\frac{C_{\ell_{i},\ell_{j}}}{n}\right)\,\,.$ (13)
The vector $\bm{\ell}$ contains the labels and gives a community structure in the case in which $C_{a,a}>C_{a,b}$ for $b\neq a$, meaning that there is a higher probability that two nodes in the same community will get connected. The value $\theta_{i}$ is proportional to the expected degree of node $i$. For this reason, if one chooses $\bm{\theta}=\mathbf{1}_{n}$ and $\bm{\ell}=\mathbf{1}_{n}$, one gets the Erdős-Renyi model, in which every node has the same expected degree, and there are no communities. The configuration model, instead, is obtained by letting $\bm{\ell}=\mathbf{1}_{n}$, but changing the value of $\bm{\theta}$ to create an arbitrary degree distribution that we choose to be a (properly rescaled) uniform distribution between $3$ and $10$ raised to the power $4$. Finally, the stochastic block model is obtained from a labeling vector different from $\mathbf{1}_{n}$ and letting $\bm{\theta}=\mathbf{1}_{n}$. We consider $k=5$ communities of equal size with $C_{a,b}=20\delta_{ab}+(1-\delta_{ab})$, with $\delta$ the Kroeneker symbol.
@@
Let $\mathcal{V}$ be a set of $n$ nodes. For each $i\in\mathcal{V}$ let $\mathbf{x}_{i}\in\mathbb{R}^{2}$ be a random vector with norm $\|\mathbf{x}_{i}\|\leq 1$. The entries of the graph adjacency matrix $A\in[0,1]^{n\times n}$ are generated independently (up to symmetry) at random with probability
$\mathbb{P}(A_{ij}=1)=e^{-\beta\|\mathbf{x}_{i}-\mathbf{x}_{j}\|}\,\,,$ (14)
@@
Note that even though the entries of $A$ are drawn at random, this model can generate graphs with a high clustering coefficient because the probability depends on the relative distance between fixed embedding vectors. In our simulations, we set $\beta=20$.
Data availability
All data used in this article are publicly available at http://www.sociopatterns.org/datasets/ and at Ref. [44].
Code availability
Commented Python code to reproduce our results is available at Ref. [44].